OpenStudy (anonymous):

how do you get the answer to square root 5x^7 square root 10x^6

OpenStudy (kc_kennylau):

\[\sqrt{5x^7}\sqrt{10x^6}\]

OpenStudy (kc_kennylau):

Start by combining the two square roots

OpenStudy (anonymous):

like terms

OpenStudy (anonymous):

what answer did yall get

OpenStudy (anonymous):

@kc_kennylau what does the x stand for

OpenStudy (anonymous):

1

OpenStudy (anonymous):

im not really sure this is confusing to me

OpenStudy (anonymous):

im not sure what to do

OpenStudy (kc_kennylau):

The x is just a variable

OpenStudy (kc_kennylau):

\[\sqrt{5x^7}\sqrt{10x^6}\\=\sqrt{50x^{13}}\\=\sqrt{2\times25x^{12}x}\\=5x^6\sqrt{2x}\]